Grade 12th passPhysical Chemistry

25.5g of h2o2 solution on decomposition gave 1.68l of o2 at stp. the percentage strength by weight of the solution is

Murali Krishna
8 Years agoGrade 12th pass

1 Answer

Sareeka
8 Years ago
2h202=2h20+o22moles=2moles+ 1mole1mole of o2 =22.4l1.68litre of o2=1.68÷22.4=0.0750.075×2=0.15mol of h2o2Molar mass of h2o2 =34 g/mol0.15×34=5.1100÷25.5×5.1=20%